John deposit 5,000 pesos in a bank paying 12% compounded quarterly. After 4 years, John decided to close her account. How much money would she be able to get from bank? ​


1
Expert's answer
2022-02-01T10:20:10-0500

Compound interest


"A=P(1+\dfrac{r}{n})^{nt}"

Given "P=5000, r=0.12, n=4, t=4"


"A=5000(1+\dfrac{0.12}{4})^{4(4)}=8023.53"

John would be able to get 8023.53 pesos from bank.